OpenStudy (anonymous):

Is it possible for a triangle to have sides with lengths: 5cm, 7cm and 10cm? Explain your answer.

OpenStudy (anonymous):

5, 7, 10 Yes, Obtuse

OpenStudy (anonymous):

yes 5+7=12>10 10+7=17>5 5+10=15>7 the sum of two sides must be greater than the third the side